# High stakes exams inflate gender gap thus leading to systematic grading errors in introductory physics

### v1

@@ Abstract
Previous research has suggested that changing the percentage of the course grade associated with exam grades in STEM courses can change the gender gap in the course. It’s also been shown that assessments with the highest stakes have the lowest (relative) scores for female students. Previous research by the authors has shown that the implementation of retake exams can eliminate the gender gap in introductory physics courses. This paper explores several different hypotheses for why retake exams are associated with a zeroed gender gap. Analyzing data from exams with different stakes, we argue that the entire gender gap on introductory physics exams may be due to the stakes associated with those exams. In other words, the data support the idea that a gender grade-gap on exams is not measuring a gender difference in the physics knowledge or physics ability of these students.
@@ I Introduction
I.1 Gender gap in physics
Gender equity gaps in introductory physics courses and exams have been studied extensively over the past several decades. Many different explanations for their existence have been proposed and investigated. Some of the leading theories for the gender gaps in physics are that women on average i) have poorer physics and/or math preparation [4, 5], ii) have less overall interest in physics [6], and/or iii) suffer more from test anxiety, which is amplified on high-stakes exams [7, 8, 9, 5, 10, 11, 12]. It’s also been shown that women don’t feel like they belong in physics as much as their male peers [13], that they have less of a science or physics identity [14], and/or have less self-efficacy in physics [5].
@@ I.1 Gender gap in physics
Exploring one potentially fruitful mitigation, Cotner and Ballen [2] show that exam gender gaps in some biological science courses could be changed by lowering the stakes associated with exams. They compared overall success in different biology courses by categorizing them as either low-stakes (less than 40% of the course grade determined by quizzes and exams) or high stakes (more than 40% of the course grade determined by quizzes and exams). The men outperformed the women in the high-stakes courses, while there was no statistical difference between the genders in low-stakes courses.
@@
Cotner et al. later suggest [7] that the reason for the change in performance when the stakes change is because women are more likely to suffer from anxiety which is amplified in high-stakes situations. In fact the connection between high-stakes exams and anxiety in women is well studied [7, 8, 9, 5, 10, 11, 12]. Maespina & Singh [5] found that women under-performed on high-stakes exams in calculus-based introductory physics courses. However, when they controlled for test anxiety, the gap became non-significant. There is evidence that test anxiety is an issue for women in other STEM majors as well. Espinosa et al. [10] found that in a large introductory biology course, women expressed higher test anxiety than men and that anxiety was a significant predictor for exam grades even when controlling for overall GPA. This is important not only in biology courses but also because a large population of introductory physics students are biology students. Similarly, Malesina et al. [9] find the same is true for bioscience students enrolled in a physics course. Salehi et al [8] study performance across three different assessment types, exams, non-exams and labs, in five different lower division STEM courses. They found that while there was a significant negative gap for women on exams, women consistently outperformed men on the other two types of assessments. Interestingly, they found that exam performance for women was negatively influenced by test anxiety in both lower and upper division courses, even though by the upper division courses the gender gap on exam scores had disappeared in both biology courses and other STEM courses.
@@
While the evidence for anxiety makes it a compelling explanation for the gender gap, we instead propose that the high-stakes exams are the cause of the gender gap, and therefore it is the course structure that deserves mitigation. As an example for why we take this approach, consider a study performed by Simmons & Heckler [16] where they simulated new course grades for students (after a course had already been completed) by reducing the weight of the exams in the overall course grade. They found that making the exams worth less of the overall course grade, equity gaps in the course were reduced for women and other historically marginalized students compared to their non-marginalized peers. In this example, women’s anxiety was not changed, but instead the course structure was changed to successfully address the inequity. While this may seem like a semantics argument, positioning the stakes as the possible cause of the gender gap (instead of anxiety) focuses the research and intervention efforts on the changeable aspects of the course, rather than the (potentially unchangeable) aspects of entire populations. Similarly, it removes the urge to “fix” any perceived inadequacies among an underrepresented group (in this case, women).
@@ I.2 Retake Exams in Physics
One possible way to lower the stakes of an exam, while still using exams as an assessment is to offer a retake option for each exam. Thus far, there is very little research on retake exams in large introductory STEM classes. Davis et al. [18] find that students in a large introductory human physiology course who took advantage of retake exams on average increase their grade by 6%. Kortemeyer et al. [19] share the benefits of retakes as learning tools in introductory physics. However, none of the examples we found examined the relationship between the retakes and the gender gap in physics courses aside from our own previous work on this same dataset.
@@ I.3 Models of achievement and assessment equity
This work, like our prior research [1] utilizes two theoretical models. We define a course as being equitable if student grades are not predictable by demographic, meaning that all demographic groups achieve the same overall grades on average. This is the definition of an Equity of Parity model of equity [20]. We also use Cotner & Ballen’s [2] Course Deficit Model. A Course Deficit Model assumes that any gaps in achievement within the course between any demographic groups are a product of the course structure. Importantly, this does not mean that inequities prior to the course do not exist, the course deficit model merely assumes that the equity gaps measured within the course are a product of the course itself. Our previous work [1] shows that by changing course structures it is possible to eliminate equity gaps and thus achieve equity of parity.
@@ II Research questions
In this paper we make the assumption that because the gender gap decreased or was entirely eliminated in courses that use a retake exam option, that the reason for increased gender gaps in the intro physics courses in our data set is likely due to the assessment strategy in courses. This leads us to test a few different hypotheses:
@@
    Does the gender gap decrease in retake courses because students identifying as women are more likely to take advantage of the retake opportunity?
@@
    Does the gender gap decrease in retake courses because students identifying as women are doing better on their second try? (Presumably because the second try has allowed them to ‘catch up’.)
@@ III Database Details
All together the database included 26,766 students with 16,485 identified as female students and 10,246 identified as male students. The exam data include 5,567 students appearing only once, 7,324 students appearing in only two of the classes and only once in each, and 13,089 students appearing exactly once in each of the three classes in the 3-course series. The three retake classes had a total of 397 students.
The number of individual exams given during the term varied from a minimum of 2 midterms given in each of seven classes to a maximum of 9 weekly exams given in each of 6 classes with the most common number given being 4 in-term exams in each of 111 classes (three of those are the retake classes). In addition, the databases from 203 classes also had final exam grades and so those classes had all exams accounted for.
@@ III.1 Classifying the Stakes of an Assessment
For the purposes of this paper we have decided to classify the ‘stakes’ of an exam into three different categories. We use the same definition of high-stakes assessments as those typically defined in the literature. High-stakes assessments are timed exams that are usually in-person and allow limited use of resources. For example, a student might be able to bring a note card, but they would not be able to use the internet during the exam. Usually these assessments are compared to low-stakes assessments like labs and/or homework. In this paper we use a different definition of low-stakes. We consider the first try on an in-term exam that has a retake option to be a ‘low-stakes’ assessment because a second try is free to any student and could completely erase that first score. We also introduce the idea of a ‘medium stakes’ assessment. We consider the exams given as the term progresses to be ‘medium stakes’ both because each individual in-term exam counts for a much smaller portion of the course grade than the final exam and because usually the class syllabus described in-course grading rules telling students that the lowest in-term exam would be dropped before calculating the course grade. These definitions are outlined in Table 1.
Table 1: Descriptions of the magnitude of the stakes associated with an exam. The stakes associated with one exam can be thought of as determined by the minimum amount that the exams can possibly contribute to the course grade divided by the total number of exams
@@
grade by full
set of exams
@@
or two midterm exams
@@ IV.1 In-term exams in retake classes
Our expectation was that students who had high scores on their first attempt would not risk lowering that score and so would not take a retake exam. We found that an average first attempt was followed by 34% of the students taking the retake a week later. The retake classes were anomalous in having zero gender gap in course grades (i.e. female students had the same average course grades as male students unlike in the regular courses) suggesting the following possible gap-closing scenario regarding retakes: i) female students may be more likely to take retake exams to improve their grade and/or ii) female students may make larger gains on retake exams. Examination of the data shows that the first possibility is wrong. We find that 78% $\pm$ 4% of male students retook at least one exam and 74% $\pm$ 3% of female students retook at least one exam. All together male students retook an average of 1.41 $\pm$ 0.09 exams and female students retook an average of 1.33 $\pm$ 0.07 exams. It seems that the two genders had about the same need for retake exams.
@@
Figure 1 shows these gender gaps for each first attempt and also after each retake exam. These results are plotted as a function of the week the exam was taken through the 10-week term. The overall averages, after all four in-term first-attempts and their retakes were graded, are shown to the right of the end of the term. Notice that students identifying as female outscored students identifying as male on three of the four first-attempt in-term exams. And the average of all first-attempt exams is slightly positive (female students out-performing male students) but the average of all in-term exams after retakes are scored is slightly negative (retake exam scores benefited male students slightly more than female students). So the idea that the retakes allowed female students to improve their first attempt scores in a way that reduced the gender gap is also untrue. This result suggests a third possible effect of allowing retake exams. iii) The knowledge that retakes are allowed lowers the stakes of each first-attempt from what it would be without retakes. So it may be that the scores are gender neutral because the stakes of the first-attempt exams are low. Specifically, we suggest that the first-attempt has lower stakes in a retake class than it does in the regular courses where the first attempt is also the last attempt. In Figure 1 we have also included the week-by-week GPA-controlled in-term exam gender gaps for three sets of non-retake classes for which the week-by-week exam pattern through the term is known. In the rest of the paper we use the ideas of low-stakes, medium-stakes, and high-stakes exams in comparing the retake classes with the 254 non-retake classes for which we have exam data.
Figure 1: As explained in the text, each in-term exam was z-scored on a class-by-class basis and the gender gap for each group of classes was determined with HLM while controlling for each student’s incoming GPA. This in-term exam score gender gap is plotted as a function of time through the 10-week term at the times the exams and retakes were taken. On the right, after the end of the term, are the overall averages for those first-attempt exams and for the recorded scores after retakes were completed. One sees that the female students generally had higher exam scores (i.e. a positive gender gap) after the first attempt and that, overall, retakes favored the male students. For reference the gender gaps for 179 non-retake classes are also shown on a week-by-week basis through their terms.
@@ IV.2 Low-stakes versus medium-stakes assessments
Instructors in this series of classes determined course grades almost completely by averaging the exam grades using a weighting of their choice. A weighting that was common (almost universal) was to 1) drop the lowest score among the in-term exams (either one in-term exam or one-half of one midterm) and then 2) weigh the final exam somewhere between 40% and 60% depending on whether the in-term exam average was higher or lower, respectively, than the final exam score. The most common numbers of in-term exams were four and eight so the grade weight of a single in-term exam ranged from roughly 5% to 15% of the course grade. We consider these exams to be medium-stakes assessments. The final exam carries several times the weight of any particular in-term exam so we consider it to be a high stakes assessment. For the retake courses we consider the first-attempt of every exam set that students take to have low stakes because the student knows that they are free to take another exam and discount the first one if they choose to.
We are interested in comparing the gender gap resulting from low-stakes in-term exams with that from medium-stakes in-term exams. For this reason we compare the gender gap in the first-attempt exams from the retake classes with the gender gaps in the 254 non-retake classes in our dataset. Since different classes had different numbers of in-term exams we will just average the z-scored in-term-exam scores for the entire term for each student in each class. Also remember that the gender gap for in-term exams in a class has a small but clear dependence on the gender gap for GPAs in that class so we will control for GPA in our analyses.
We model the comparison of the 3 retake classes with the 254 regular classes using the same kind of 3-level hierarchical regression analysis that was previously discussed. Again, the student level is the lowest level, then students are grouped in classes (2nd level), and the classes are grouped within the course series (A, B, and C in the series identify the highest level). We use all three levels because our database covers all three levels and also because the different classes may have slightly different grade penalties for female students. We again use the variables $GPA$ for a student’s GPA, $Female$ = 1 for female students and = 0 for male students, and now include $Retake$ = 1 for the three retake classes and = 0 for the 254 regular classes. The model we fit for the in-term exam z-scores is
@@
where the coefficient $b_{Female}$ represents the GPA-controlled gender gap and the coefficient $b_{Female*Retake}$ measures the difference in that gender gap when retakes are possible.
@@
Note that even though the coefficient for $Retake$ is negative, indicating (perhaps) that men do worse under retake conditions, this is only because the grades are z-scored. In our previous paper [1] we showed that, on average, both genders had higher grades in the retake classes. Also, as shown in Appendix C, the gender gap decreases in size in retake classses for each racial/ethnic group which are sufficiently represented in our data. Finally, as discussed in Appendix B, to decide if these coefficients or our resulting conclusions are affected by the fact that many students are in our database more than once (because they have taken more than one of the classes in this three-term series) we have tested the effects of limiting each student to just one appearance in the database. We find that the coefficients and our resulting conclusions are unchanged.

### v2

# High stakes exams inflate a gender gap and contribute to systematic grading errors in introductory physics

@@ Abstract
Previous research has suggested that changing the percentage of the course grade associated with exam grades in STEM courses can change the gender gap in the course. It has also been shown that assessments with the highest stakes have the lowest (relative) scores for female students. Previous research by the authors has shown that the implementation of retake exams can eliminate the gender gap in introductory physics courses. This paper explores several different hypotheses for why retake exams are associated with a zeroed gender gap. Two independent measurements comparing exams with different stakes are used in support of the argument that the entire gender gap on introductory physics exams may be due to the stakes associated with those exams. In other words, these data support the idea that a gender grade gap on exams is not measuring a gender difference in the physics knowledge or physics ability of these students. Implications suggest that instructors should choose lower stakes assessment options if they are interested in exam measurements that are not influenced by differences in students’ performance related to exam stakes.
@@ I Introduction
I.1 Gender gaps in physics
Gender equity gaps in introductory physics courses and exams have been studied extensively over the past several decades. Gender gaps in physics can refer to performance gaps in course grades [4, 1], on concept inventories [5, 6], and/or on exams [7]. Representation and retention gaps have also been observed [8].
I.1.1 Gaps - student level characteristics
Many different explanations for the existence of these types of gender gaps have been proposed and investigated. Some of the leading theories for the gender gaps in physics are that women on average i) have poorer physics and/or math preparation [9, 10, 6], ii) have less overall interest in physics [11], and/or iii) suffer more from test anxiety, which is amplified on high-stakes exams [12, 13, 14, 10, 15, 16, 7]. It’s also been shown that women don’t feel like they belong in physics as much as their male peers [17, 4], that they have less of a science or physics identity [18], and/or have lower self-efficacy in physics [10, 15, 19, 20].
@@ I.1.1 Gaps - student level characteristics
Exploring one potentially fruitful mitigation of gender gaps in exams, Cotner and Ballen [2] show that exam gender gaps in some biological science courses could be changed by lowering the stakes associated with exams. They compared overall success in different biology courses by categorizing them as either low-stakes (less than 40% of the course grade determined by quizzes and exams) or high stakes (more than 40% of the course grade determined by quizzes and exams). The men outperformed the women in the high-stakes courses, while there was no statistical difference between the genders in low-stakes courses.
@@
Cotner et al. later suggest [12] that the reason for the change in performance when the stakes change is because women are more likely to suffer from anxiety which is amplified in high-stakes situations. In fact the connection between high-stakes exams and anxiety in women is well studied [12, 13, 14, 10, 15, 16, 7], where evidence suggests that women are more likely to suffer from test-anxiety that negatively impacts exam performance [21]. Malespina & Singh [10] found that women under-performed on high-stakes exams in calculus-based introductory physics courses. However, when they controlled for test anxiety, the gap became non-significant. There is evidence that test anxiety is an issue for women in other STEM majors as well. Espinosa et al. [15] found that in a large introductory biology course, women expressed higher test anxiety than men and that anxiety was a significant predictor for exam grades even when controlling for overall GPA. This is important not only in biology courses but also because a large population of introductory physics students are biology students. Similarly, Malespina et al. [14] find the same is true for bioscience students enrolled in a physics course. Salehi et al [13] study performance across three different assessment types, exams, non-exams and labs, in five different lower division STEM courses. They found that while there was a significant negative gap for women on exams, women consistently outperformed men on the other two types of assessments. Interestingly, they found that exam performance for women was negatively influenced by test anxiety in both lower and upper division courses, even though by the upper division courses the gender gap on exam scores had disappeared in both biology courses and other STEM courses.
I.1.2 Gaps - system level issues
Leslie et al. [22] take a more systemic view in noting that gender gaps in representation are largest in a few specific academic fields (including physics) where the practitioners in the field are more likely to think that to succeed in their field “hard work alone just won’t cut it; you need to have an innate gift or talent.” Muradoglu et al. [23] use this idea, that success in a field requires innate brilliance, together with the general societal view that women are not as intellectually talented as men [24], to argue that a natural result is that, on average, women would have lower average self-efficacy in such fields. Since, self-efficacy inversely correlates with anxiety, it also is reasonable that women, on average, would be expected to be more anxious than men during the high stakes exams in these particular academic fields. As noted earlier, in this paper we will show data consistent with the idea that there is no gender gap in physics understanding/ability but that the commonly measured exam gender gap arises because of the stakes associated with that exam. In this sense the exam gender gap may simply be an incorrect measurement of the underlying knowledge/ability. It would be ironic if such illusory exam gender grade gaps were part of the evidence society used to conclude that women were not as talented as men at physics and that that conclusion led to women being more anxious on, and under-performing on, exactly these high stakes exams that led to society’s conclusions.
I.1.3 Changes in course structure changes gaps
Although the evidence for anxiety may make it a compelling explanation for the gender gap, we will continue our previous argument [1] that it is the course structure that determines demographic grade gaps. We suggest that high-stakes exams be considered the cause of the gender gap. One reason we take this position is because high-stakes exams (being characteristics of the course structure) are easily changed, while anxiety (being a characteristic of the student) is much more difficult to change. If high stakes exams are the cause of the gender gap, changes in course structure can serve as a mitigation.
@@ I.1.3 Changes in course structure changes gaps
As an example for why we take this approach, consider a study performed by Simmons & Heckler [25] where they simulated new course grades for students (after a course had already been completed) by reducing the weight of the exams in the overall course grade. They found that making the exams worth less of the overall course grade, equity gaps in the course were reduced for women and other historically marginalized students compared to their non-marginalized peers. In this example, women’s anxiety was not changed, but instead the course structure was changed to successfully address the inequity. While this may seem like a semantics argument, positioning the stakes as the possible cause of the gender gap (instead of anxiety) focuses the research and intervention efforts on the changeable aspects of the course, rather than the (potentially unchangeable) aspects of entire populations. Similarly, it removes the urge to “fix” any perceived inadequacies among an underrepresented group (in this case, women).
@@ I.2 Retake Exams in Physics
One possible way to lower the stakes of an exam, while still using exams as an assessment, is to offer a retake option for each exam given during the term. For each in-term exam, a retake covers the same topics as the first-attempt exam, is offered after the students have seen their score on their first-attempt, and is given a score that can supplant the first-attempt score. Thus far, there is very little research on retake exams in large introductory STEM classes. Davis et al. [26] find that students in a large introductory human physiology course who took advantage of retake exams on average increase their grade by 6%. Kortemeyer et al. [27] share the benefits of retakes as learning tools in introductory physics. However, none of the examples we found examined the relationship between the retakes and the gender gap in physics courses aside from our own previous work on this same dataset.
@@ I.3 Models of achievement and assessment equity
This work, like our prior research [1] utilizes two theoretical models. We define a course as being equitable if student grades are not predictable by demographic, meaning that all demographic groups achieve the same overall grades on average. This is the definition of an Equity of Parity model of equity [28]. We also use Cotner & Ballen’s [2] Course Deficit Model. A Course Deficit Model assumes that any gaps in achievement within the course between any demographic groups are a product of the course structure. Importantly, this does not mean that inequities prior to the course do not exist, the course deficit model merely assumes that the equity gaps measured within the course are a product of the course itself. Our previous work [1] shows that by changing course structures it is possible to eliminate equity gaps and thus achieve equity of parity without changing the course content or the intellectual level of the content.
@@ II Research questions
In this paper we make the assumption that because the gender gap decreased or was entirely eliminated in courses that use a retake exam option, that the reason for increased gender gaps in the intro physics courses in our data set is likely due to the assessment strategy in courses. This leads us to a few different research questions:
@@
    Does the gender gap decrease, in course sections that allow retakes on exams, because women students are more likely to take advantage of the retake opportunity?
@@
    Does the gender gap decrease, in course sections that allow retakes on exams, because women students are doing better on their second try?
@@
Each research question tests an assumption that either came from the literature or was proposed to the authors of this paper in conversations about our previous work [1]. The first research question examines the assumption that women and men have behavioral differences that influence how likely they are to take advantage of optional opportunities to improve their grade. The second research question tests the assumption that because women are less likely to be prepared to take physics, that they might need more time to practice and receive help before succeeding than men do. The third research question was born out of negative results from the first two questions.
@@ III Database Details
All together the database included 26,783 students with 16,493 identified as female students, 10,254 identified as male students, 5 identified as non-binary, and 31 students with unknown gender identities. This university started collecting self-identified gender information in the Fall term of 2015 so our dataset is almost completely binary gender only. For completeness, the data regarding the small numbers of unknown gender and non-binary students are discussed in Appendix F. The exam data include 5,598 students appearing only once, 7,382 students appearing in only two of the classes and only once in each, and 13,268 students appearing exactly once in each of the three classes in the 3-course series. The three retake classes had a total of 397 students.
The number of individual exams given during the term varied from a minimum of 2 midterms given in each of seven classes to a maximum of 9 weekly exams given in each of 6 classes with the most common number given being 4 in-term exams in each of 115 classes (three of those are the retake classes). In addition, the databases from 203 classes also had final exam grades and so those classes had all exams accounted for.
@@ III.1 Classifying the Stakes of an Assessment
For the purposes of this paper we will classify the stakes of an exam into three different categories. We use the same definition of high-stakes assessments as those typically used in the literature [2]. “High-stakes” assessments are timed exams that are usually in-person and allow limited use of resources. For example, a student might be able to bring a note card, but they would not be able to use the internet during the exam. In addition, these exams are considered high-stakes because their scores are a large factor in determining the course grade. This is an implied distinction in previous work because high-stakes exams are often differentiated from quizzes that are also usually timed and allow limited resources [2] but have scores that are not as important factors in the course grade.
Instead of examining stakes of different assessment types (e.g. homework vs. exams) as other have done, in this paper we are investigating if the differences in the stakes of different exams have an impact on equity gaps found in those exams. To that end, we will take the stakes of any particular exam to be a measure of the importance of that exam’s (or quiz’s) score to the course grade. For instance, in the classes included in our database, the grade on the final exam constitutes roughly 50% of the course grade (no courses in our dataset allowed retakes on final exams). Therefore, the final exams in either type of course are considered “high-stakes” as compared to other exams.
We also introduce the idea of a medium stakes assessment. In non-retake courses, the entire set of two to nine in-term exams (in-term meaning excluding the final) constitute roughly the other 50% of the course grade so each of those exams has considerably lower stakes than the lone final exam. We consider these exams given as the term progresses to be “medium stakes” both because each individual in-term exam counts for a much smaller portion of the course grade than the final exam and because usually the class syllabus described in-course grading rules telling students that the lowest in-term exam would be dropped before calculating the course grade.
On the other hand, in retake courses, the entire set of four first-attempt in-term exam scores may have no contribution at all to the course grade because the retake exam score is always available to supplant the first-attempt exam score. For this reason we consider a first-attempt exam in a retake course to have relatively lower stakes than any of the in-term exams in a non-retake class, and therefore define the first attempt on a retake as a “low-stakes” assessment.
We consider the first try on an in-term exam that has a retake option to be a low-stakes assessment because a second try is free to any student and could completely erase that first score. But we consider the second attempt in a retake class to be a medium stakes assessment because this grade will contribute to the final grade at roughly the same weight of the in-term exams in the courses that do not have retake exams. All of these definitions of the stakes of different exams are outlined in Table 1.
Note that the medium-stakes exams in the retake classes (the actual retake exams) are only taken by a subset of students who choose to take advantage of the opportunity to do so. This will make it difficult to draw conclusions from this data set so we will leave any calculations of gender gaps within these student subsets, varying through the term, to Appendix E.
Table 1: Descriptions of the magnitude of the stakes associated with an exam. The stakes associated with one exam can be thought of as determined by the minimum amount that that set of exams can possibly contribute to the course grade divided by the total number of exams of that type. Retake courses and regular courses each had a single final exam so all final exams had high stakes.
@@
grade by an
average exam
@@
Medium or two midterm exams 5%-15%
Retake class - Second-
attempt at in-term
exams (i.e. the retakes)
@@ IV.1 In-term exams in retake classes
Our expectation was that students who had high scores on their first attempt would not risk lowering that score and so would not take a retake exam. We found that on average a first attempt of an exam was followed by 34% of the students taking the retake a week later. The retake classes were anomalous in having zero gender gap in course grades (i.e. female students had the same average course grades as male students unlike in the regular courses) suggesting the following possible gap-closing scenario regarding retakes: i) female students may be more likely to take retake exams to improve their grade and/or ii) female students may make larger gains on retake exams. Examination of the data shows that the first possibility is wrong. We find that 78% $\pm$ 4% of male students retook at least one exam and 74% $\pm$ 3% of female students retook at least one exam. All together male students retook an average of 1.41 $\pm$ 0.09 exams and female students retook an average of 1.33 $\pm$ 0.07 exams. It seems that the two genders had about the same need for retake exams.
@@
Figure 1 shows these gender gaps for each first attempt and also after each retake exam. These results are plotted as a function of the week the exam was taken through the 10-week term. The overall averages, after all four in-term first-attempts and their retakes were graded, are shown to the right of the end of the term. Notice that female students outscored male students on three of the four first-attempt in-term exams. And the average of all first-attempt exams is slightly positive (female students out-performing male students) but the average of all in-term exams after retakes are scored is slightly negative (retake exam scores benefited male students slightly more than female students). So the idea that the retakes allowed female students to improve their first attempt scores in a way that reduced the gender gap is also untrue. This result suggests a third possible effect of allowing retake exams. iii) The knowledge that retakes are allowed lowers the stakes of each first-attempt from what it would be without retakes. So it may be that the scores are gender neutral because the stakes of the first-attempt exams are low. Specifically, we suggest that the first-attempt has lower stakes in a retake class than it does in the regular nonretake classes where the first attempt is also the last attempt. In Figure 1 we have also included the week-by-week GPA-controlled in-term exam gender gaps for three sets of non-retake classes for which the week-by-week exam pattern through the term is known. In the rest of the paper we use the ideas of low-stakes, medium-stakes, and high-stakes exams in comparing the retake classes with the 254 non-retake classes for which we have exam data.
Figure 1: Each in-term exam was normalized on a class-by-class basis so that the standard deviation of the grade distribution is one. The gender gap for each group of classes was determined with Hierarchical Linear Modeling while controlling for each student’s incoming GPA. This in-term exam score gender gap is plotted as a function of time through the 10-week term at the times the exams and retakes were taken. On the right, after the end of the term, are the overall averages for those first-attempt exams and for the recorded scores after retakes were completed. One sees that the female students generally had higher exam scores than male students (i.e. a positive gender gap) after the first attempt and that, overall, retakes favored the male students. For reference the gender gaps for 180 non-retake classes are also shown on a week-by-week basis through their terms.
@@ IV.2 Low-stakes versus medium-stakes assessments
Instructors in this series of classes determined course grades almost completely by averaging the exam grades using a weighting of their choice. A weighting that was common (almost universal) was to 1) drop the lowest score among the in-term exams (either one in-term exam or one-half of one midterm) and then 2) weigh the final exam somewhere between 40% and 60% depending on whether the in-term exam average was higher or lower, respectively, than the final exam score. The most common numbers of in-term exams were four and eight so the grade weight of a single in-term exam ranged from roughly 5% to 15% of the course grade. We consider these exams to be medium-stakes assessments. The final exam carries several times the weight of any particular in-term exam so we consider it to be a high stakes assessment. For the retake courses we consider the first-attempt of every exam set that students take to have low stakes because the student knows that they are free to take another exam and discount the first one if they choose to. In all classes (retake and non-retake) it was not possible to retake the final exam.
We are interested in comparing the gender gap resulting from low-stakes in-term exams with that from medium-stakes in-term exams. For this reason we compare the gender gap in the first-attempt exams from the retake classes with the gender gaps in the 254 non-retake classes in our dataset. Since different classes had different numbers of in-term exams we will just average the z-scored in-term-exam scores for the entire term for each student in each class and model the term-averaged in-term exam z-score for each student. Also remember that the gender gap for in-term exams in a class has a small but clear dependence on the gender gap for GPAs in that class so we will control for GPA in our analyses.
We model the comparison of the 3 retake classes with the 254 regular classes using the same kind of 3-level hierarchical regression analysis that was previously discussed. Again, the student level is the lowest level, then students are grouped in classes (2nd level), and the classes are grouped within the course series (A, B, and C in the series identify the highest level). We use all three levels because our database covers all three levels and also because the different classes may have slightly different grade penalties for female students. We again use the variables $GPA$ for a student’s GPA, $Female$ = 1 for female students and = 0 for male students, and now include $Retake$ = 1 for the three retake classes and = 0 for the 254 regular classes. The model we fit for the student’s in-term exam z-score is
@@
where the coefficient $b_{Female}$ represents the GPA-controlled gender gap and the coefficient $b_{Female*Retake}$ measures the change in the gender gap when retakes are possible.
@@
Note that even though the coefficient for $Retake$ is negative, indicating (perhaps) that men do worse under retake conditions, this is only because the grades are z-scored. In our previous paper [1] we showed that, on average, both genders had higher grades in the retake classes. Also, as shown in Appendix C, the gender gap decreases in size in retake classes for each racial/ethnic group which are sufficiently represented in our data. Finally, as discussed in Appendix B, to decide if these coefficients or our resulting conclusions are affected by the fact that many students are in our database more than once (because they have taken more than one of the classes in this three-term series) we have tested the effects of limiting each student to just one appearance in the database. We find that the coefficients and our resulting conclusions are unchanged.